OpenStudy (anonymous):

Two groups of students order burritos and tacos at a local restaurant. One order of 3 burritos and 4 tacos costs $11.33. The other order of 9 burritos and 5 tacos costs $23.56. What is the cost of a taco?

OpenStudy (texaschic101):

3b + 4t = 11.33 --->(-3) 9b + 5t = 23.56 --------------- -9b - 12t = -33.99 (result of multiplying by -3) 9b + 5t = 23.56 ---------------add - 7t = - 10.43 t = -10.43/-7 t = 1.49 I am going to complete the problem just to check the answer now sub 1.49 in for t in either equation to find b 3b + 4t = 11.33 3b + 4(1.49) = 11.33 3b + 5.96 = 11.33 3b = 11.33 - 5.96 3b = 5.37 b = 5.37/3 b = 1.79 check.. 9b + 5t = 23.56 9(1.79) + 5(1.49) = 23.56 16.11 + 7.45 = 23.56 23.56 = 23.56 (correct) burritos (b) cost $ 1.79 tacos (t) cost $ 1.49